Please help. A plane made a flight with the wind between 2 towns in 2 hours. Returning against the wind, the planes speed was 60 mph slower and the flight took 3 hours. What was the planes speed in each direction.
Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

2X=3(X-60) OR 2X=3X-180 OR X=180 OR WHICH IS THE SPEED FOR THE 2 HOUR TRIP OR 2*180=360
& 3(X-60) OR 3(180-60) OR 3(120)=360

Answer by stanbon(75887)

A plane made a flight with the wind between 2 towns in 2 hours. Returning against the wind, the planes speed was 60 mph slower and the flight took 3 hours. What was the planes speed in each direction.
With the wind DATA:
distance = d; rate=r ; time= d/r = d/r=2 hrs.
Against the wind DATA:
distance = d; rate=r-60; time=d/r = d/(r-60)=3 hrs.
EQUATION:
d=2r
d=3(r-60)
Therefore 2r=3(r-60)
2r=3r-180
-r=-180
r=180 mph (with the wind rate)
r-60=120 mph (against the wind rate)
Cheers,
Stan H.